what are the factors of (p-q)3+(q-r)3+(r-p)3 helpme with explanation

Grade:8

1 Answers

Vijay Mukati
8 years ago
Dear Student,
If a+b+c = 0,....(1)
then a3 + b3 + c3 = 3abc.
In your cas a = p-q, b = q-r and c = r-p.
If it is also satisfying the (1) condition.
Therefore the factors will be 3abc = 3(p-q)(q-r)(r-p).
Thanks.